# Boosting Adversarial Transferability with Low-Cost Optimization via Maximin Expected Flatness

### v2

@@ Abstract
Transfer-based attacks craft adversarial examples on white-box surrogate models and directly deploy them against black-box target models, offering model-agnostic and query-free threat scenarios. While flatness-enhanced methods have recently emerged to improve transferability by enhancing the loss surface flatness of adversarial examples, their divergent flatness definitions and heuristic attack designs suffer from unexamined optimization limitations and missing theoretical foundation, thus constraining their effectiveness and efficiency. This work exposes the severely imbalanced exploitation-exploration dynamics in flatness optimization, establishing the first theoretical foundation for flatness-based transferability and proposing a principled framework to overcome these optimization pitfalls. Specifically, we systematically unify fragmented flatness definitions across existing methods, revealing their imbalanced optimization limitations in over-exploration of sensitivity peaks or over-exploitation of local plateaus. To resolve these issues, we rigorously formalize average-case flatness and transferability gaps, proving that enhancing zeroth-order average-case flatness minimizes cross-model discrepancies. Building on this theory, we design a Maximin Expected Flatness (MEF) attack that enhances zeroth-order average-case flatness while balancing flatness exploration and exploitation. Extensive evaluations across 22 models and 24 current transfer-based attacks demonstrate MEF’s superiority: it surpasses the state-of-the-art PGN attack by 4% in attack success rate at half the computational cost and achieves 8% higher success rate under the same budget. When combined with input augmentation, MEF attains 15% additional gains against defense-equipped models, establishing new robustness benchmarks. Our code is available at https://github.com/SignedQiu/MEFAttack.
Index Terms:
Adversarial attack, Black-box attack, Adversarial transferability, Loss surface flatness, Loss landscape
@@
Figure 1: Visualization of adversarial loss landscapes for nine attacks on Res-50 [9]. The loss surfaces is constructed by perturbing adversarial examples along two random directions. Our MEF attack achieves the flattest loss landscape, indicating superior flatness-based transferability.
@@ I Introduction
Adversarial examples [10, 11, 12, 13] pose a significant security threat to deep neural networks (DNNs). By introducing imperceptible perturbations to benign inputs, attackers can craft adversarial examples that mislead DNNs into producing erroneous predictions. Early adversarial attacks primarily operated under two paradigms: white-box attacks that require full access to the target model’s architecture and parameters, such as FGSM [14], PGD [15], and C&W [13], and query-based grey-box attacks that rely on iterative adversarial direction estimation through repeated model queries, as seen in Boundary Attack [16], SimBA [17], and HSJA [18]. While effective, these methods face critical deployment barriers: white-box attacks assume unrealistic access to victim models, while query-based methods incur prohibitive query overhead, making them impractical in real-world scenarios.
Transfer-based attacks [19, 20, 21, 1, 22, 2, 23, 24] overcome these limitations by crafting adversarial examples on surrogate models and directly transferring them to black-box targets. This paradigm eliminates both dependency on target model’s internals and query interactions, making it a practical threat to real-world systems. Among these attacks, a recent class of flatness-enhanced methods [3, 7, 8, 2] has emerged as a promising direction by linking the geometric flatness of the adversarial loss landscape to adversarial transferability. Flatness-enhanced methods craft adversarial examples with flatter neighborhood loss landscapes, enhancing cross-model robustness and achieving state-of-the-art transferability [8].
@@
Based on the empirical and theoretical analysis, we propose the Maximin Expected Flatness (MEF) attack to enhance adversarial transferability while reducing computational cost. Our key insight is to leverage the complementary strengths of average-case and worst-case flatness optimization, where average-case methods stabilize local exploitation while worst-case strategies preserve global exploration capability. MEF attack explicitly optimizes zeroth-order worst-neighborhood average flatness, a unified metric that internally computes neighborhood-level average-case flatness to stabilize local exploitation and externally identifies worst-case regions to guide global exploration. MEF implements this via gradient-guided neighborhood conditional sampling (NCS) for efficient worst-region identification and gradient balancing optimization (GBO) that reuses gradients across stages, reducing computation by 50% while enhancing transferability. By coupling these mechanisms, MEF achieves flattest loss landscape (as shown in Fig. 1) and enhanced adversarial transferability. Third-party evaluation on the widely-adopted TransferAttack benchmark shows MEF dominates 22 gradient-based methods as the new leader. Our main contributions are summarized as follows:
@@
The paper is organized as follows. Sec. II reviews transfer-based attacks and defenses. Sec. III unifies fragmented flatness definitions and reveals the limitations in current flatness optimization. Sec. IV establishes the theoretical foundation for flatness-based transferability. Sec. V introduces Maximin Expected Flatness (MEF) attack. Sec. VI conducts extensive evaluation experiments and Sec. VII concludes the paper.
@@ II Related Work
This section systematically reviews adversarial machine learning literature through two lenses: (1) we survey adversarial attacks with emphasis on transfer-based attacks, and (2) analyze various adversarial defense mechanisms, including both reactive and proactive defenses.
@@ II-A Adversarial Attacks
$\underset{\mathbf{x}^{adv}\in B_{\epsilon}(\mathbf{x})}{\max}J(\mathbf{x}^{adv},y;\mathcal{F}),$
@@
Data-driven Methods. Data-driven methods [19, 20, 22, 25] boost adversarial transferability by aggregating gradients over multiple augmented variants through data transformation strategies. The key distinction between these methods centers on the specific transformation operators. DI [19] optimizes over resized images, TI [20] applies spatial shifting, and SI [22] manipulates pixel scaling. Beyond these single-input spatial transformations, Admix [25] enhances diversity by blending cross-category images, while SSA [26] perturbs frequency domains to break spatial locality. To further amplify input variety, composite strategies integrate multiple transformations [27, 28]. However, such strategies incur significant computational overheads and induce semantic distortions [29], ultimately compromising their effectiveness.
@@
Model-driven Methods. Model-driven methods [30, 23, 31, 32, 33] exploit geometric priors [33] in surrogate models’ feature spaces to enhance adversarial transferability, primarily through two strategies: model ensemble aligns heterogeneous architectures via joint perturbation optimization to amplify spectral commonality with victims [31, 32], while surrogate refinement adversarially tunes surrogates or aligns gradients to reshape decision boundaries [30, 23]. However, model ensemble scales linearly with surrogates’ computational costs, and refinement struggles on large datasets due to complex regularization, limiting both to small-scale applications [34].
@@
Optimization-driven Methods. Optimization-driven methods [21, 7, 35, 24], enhance transferability by redesigning optimization objectives [35, 24] or algorithms [36, 21]. Advanced Objective strategies replace cross-entropy loss with feature discrepancy losses that align surrogate-victim model features [35, 24]. Generative Modeling strategies leverage generative networks to synthesize perturbations without iterative updates [36]. Gradient Stabilization strategies retain simple losses but refines gradient dynamics via momentum [21, 22, 4] or variance reduction [1]. Compared to the former two paradigms requiring complex losses or architectures, gradient stabilization strategies achieve transferability with minimal complexity, avoiding auxiliary data and models.
@@
Recent advances in gradient stabilization attacks have seen the emergence of flatness-enhanced methods, which propose that adversarial examples within flatter loss regions exhibit stronger transferability. These methods diverge into two strategies: those optimizing average-case flatness by aggregating gradients over local neighborhoods (e.g., FEM [3], APP [6]), and those pursuing worst-case flatness by minimizing gradient peaks (PGN [8]) or maximizing minimal loss (RAP [7]). However, existing works adopt fragmented flatness definitions while relying on empirically assumed correlations between flatness and transferability, obscuring theoretical understanding of their inherent limitations. Our work bridges this gap by establishing unified flatness-transferability guarantees and adaptive optimization that harmonizes both flatness criteria.
@@ II-B Adversarial Defenses
Proactive Defenses. Proactive defenses aim to intrinsically harden model parameters against adversarial perturbations. The dominant approach, adversarial training (AT) [15, 37], reformulates robust model optimization as a min-max optimization problem, where the inner maximization generates perturbations to craft challenging adversarial examples, while the outer minimization enforces the model to correctly classify these perturbed samples. Variants like ensemble adversarial training (EAT) [38] augment robustness through multi-model perturbations during training, strengthening defense against cross-architecture transfer attacks. In contrast to AT’s empirical robustness, randomized smoothing [39] provides certifiable robustness by injecting Gaussian noise into inputs and certifying predictions over perturbed samples. While this method theoretically guarantees robustness within a certified radius, such radii are often smaller than adversarial perturbations defended by AT, rendering it ineffective against practical adversarial threats despite its theoretical appeal [40].
Reactive Defenses. Reactive defenses mitigate adversarial attacks without modifying model parameters. Adversarial purification [41, 42, 43] projects inputs onto clean data manifolds using generative models [41, 43] or denoisers [42], while adversarial detection identifies adversarial examples via activation anomalies or confidence thresholds [44]. These strategies offer plug-and-play deployment but suffer from two limitations: purification efficacy depends on generative model capacity [45], and detectors can be bypassed by adaptive attacks that mimic clean data statistics [46].
Our attack demonstrates strong transferability against both proactive defenses [38, 39] and reactive defenses [41, 42, 43], effectively bypassing their protection mechanisms through flatness-enhanced perturbation crafting.
@@ III Limitations in Flatness Enhancement
The heuristic adoption of flatness concepts in flatness-enhanced methods lacks rigorous theoretical grounding while suffering from absence of systematic comparisons between flatness metrics, which constrains their effectiveness and efficiency. In this section, we first formalizes flatness as a measurable geometric property, resolving definitional ambiguities across existing methods. Then we identify two inherent limitations in current flatness-enhanced methods. Finally, we distill the critical yet understudied challenges in balancing local exploitation and global exploration for real-world deployment.
@@ III-A K-th Order Flatness
    $\widehat{R}^{(n)}_{\xi}(\mathbf{x})\;\triangleq\;\max_{\mathbf{x}^{\prime}\in B_{\xi}(\mathbf{x})}\big{\|}\nabla^{n}J(\mathbf{x}^{\prime})\;-\;\nabla^{n}J(\mathbf{x})\big{\|}_{2},$
@@
    $\overline{R}^{(n)}_{\xi}(\mathbf{x})\;\triangleq\;\mathbb{E}_{\mathbf{x}^{\prime}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}))}\big{[}\big{\|}\nabla^{n}J(\mathbf{x}^{\prime})\;-\;\nabla^{n}J(\mathbf{x})\big{\|}_{2}\big{]},$
where $B_{\xi}(\mathbf{x})=\big{\{}\mathbf{x}^{\prime}\in\mathbb{R}^{d}\,\big{|}\,\|\mathbf{x}^{\prime}-\mathbf{x}\|_{\infty}\leq\xi\big{\}}$ denotes the closed $\ell_{\infty}$‑ball centered at $\mathbf{x}$ with radius $\xi$, $\nabla^{n}J(\mathbf{x})$ denotes the $n$‑th order derivative of $J$ at $\mathbf{x}$ and $\mathrm{Unif}(B_{\xi}(\mathbf{x}))$ denotes the uniform distribution on the set $B_{\xi}(\mathbf{x})$.
In Table I, we summarize six typical flatness-enhanced methods through the lens of our formalized $n$-order flatness metrics. Current methods adopt four distinct flatness criteria: $\widehat{R}^{(0)}_{\xi}$, $\widehat{R}^{(1)}_{\xi}$, $\overline{R}^{(0)}_{\xi}$, and $\overline{R}^{(1)}_{\xi}$. While RAP [7], GNP [5], and PGN [8] optimize worst-case flatness via maximal loss ($\widehat{R}^{(0)}_{\xi}$) or gradient norm ($\widehat{R}^{(1)}_{\xi}$) constraints, FEM [3], APP [6], and TPA [2] focus on average-case objectives, either expected loss ($\overline{R}^{(0)}_{\xi}$) or gradient magnitudes ($\overline{R}^{(1)}_{\xi}$).
@@ III-B Rethinking Flatness-Enhanced Methods
The definition duality between average- ($\overline{R}^{(n)}_{\xi}$) and worst-case ($\widehat{R}^{(n)}_{\xi}$) flatness fundamentally governs their optimization behaviors. The expectation in $\overline{R}^{(n)}_{\xi}$ aggregates local gradient variations, prioritizing exploitation of dominant neighborhood. While encouraging smoother updates, this gradient averaging risks over-exploitation to suboptimal plateaus dominated by local gradients. In contrast, $\widehat{R}^{(n)}_{\xi}$’s maximization explicitly probes extreme sensitivity points, demanding exhaustive exploration to identify maximal $n$-th derivative discrepancies. Such over-exploration, however, often oscillates between sensitivity peaks, particularly under practical sampling constraints, due to the non-convexity of adversarial landscapes [15, 13].
To validate the over-exploitation and over-exploration issues, we conduct an experiment on 100 ImageNet samples [47], running four flatness-enhanced methods (RAP [7], FEM [3], PGN [8], TPA [2]) for 100 iterations. We measure inter-update gradient cosine similarity during optimization on Res-50 [9] and evaluate attack success rates against Inc-v3 [48]. The experimental results, illustrated in Fig. 2, reveal a clear divergence. For both methods, worst-case flatness optimization exhibit significantly lower inter-iteration gradient similarities (mean similarity: RAP 0.10, PGN 0.25) compared to average-case counterparts (FEM 0.38, TPA 0.88) after convergence. The persistently low gradient similarity in worst-case optimization reveals unstable oscillation between sensitivity peaks, an indicator of over-exploration. In contrast, high similarity in average-case optimization suggests over-exploitation of local gradient consensus, leading to premature convergence on suboptimal plateaus.
Crucially, both over-exploitation and over-exploration coincide with stagnated ASR improvements, confirming our prior analysis: worst-case flatness optimization exhaustively explores non-informative sensitivity extrema, while average-case criteria over-exploit locally dominant gradient directions. These inherent limitations in both flatness criteria fundamentally constrain further enhancements to both landscape flatness and adversarial transferability, establishing a critical yet understudied challenge in balancing exploitation and exploration.

### v3

@@ Abstract
Transfer-based attacks craft adversarial examples on white-box surrogate models and directly deploy them against black-box target models, offering practical query-free threat scenarios. While flatness-enhanced methods have recently emerged to improve transferability by enhancing the loss surface flatness of adversarial examples, their divergent flatness definitions and heuristic attack designs suffer from unexamined optimization limitations and missing theoretical foundation, thus constraining their effectiveness and efficiency. This work exposes the severely imbalanced exploitation-exploration dynamics in flatness optimization, establishing the first theoretical foundation for flatness-based transferability and proposing a principled framework to overcome these optimization pitfalls. Specifically, we systematically unify fragmented flatness definitions across existing methods, revealing their imbalanced optimization limitations in over-exploration of sensitivity peaks or over-exploitation of local plateaus. To resolve these issues, we rigorously formalize average-case flatness and transferability gaps, proving that enhancing zeroth-order average-case flatness minimizes cross-model discrepancies. Building on this theory, we design a Maximin Expected Flatness (MEF) attack that enhances zeroth-order average-case flatness while balancing flatness exploration and exploitation. Extensive evaluations across 33 models and 43 current transfer-based attacks demonstrate MEF’s superiority: it surpasses the state-of-the-art PGN attack by 4% in attack success rate at half the computational cost and achieves 8% higher success rate under the same budget. When combined with input augmentation, MEF attains 15% additional gains against defense-equipped models, establishing new robustness benchmarks. Our code is available at https://github.com/SignedQiu/MEFAttack.
@@
Figure 1: Visualization of adversarial loss landscapes for nine attacks on Res-50 [9]. The loss surfaces are constructed by perturbing adversarial examples along two random directions. Our MEF attack achieves the flattest loss landscape, indicating superior flatness-based transferability.
@@ I Introduction
Adversarial examples [10, 11] pose a significant security threat to deep neural networks (DNNs). By introducing imperceptible perturbations to benign inputs, attackers can craft adversarial examples that mislead DNNs into producing erroneous predictions. Early adversarial attacks primarily operated under two paradigms: white-box attacks that require full access to the target model’s architecture and parameters, such as L-BFGS [10] and PGD [11], and query-based grey-box attacks that rely on iterative adversarial direction estimation through repeated model queries, as seen in Boundary Attack [12] and SimBA [13]. While effective, these methods face critical deployment barriers: white-box attacks assume unrealistic access to victim models, while query-based methods incur prohibitive query overhead, making them impractical in real-world scenarios.
Transfer-based attacks [14, 15, 16, 1, 17, 2] overcome these limitations by crafting adversarial examples on surrogate models and directly transferring them to black-box targets. This paradigm eliminates both dependency on target model’s internals and query interactions, making it a practical threat to real-world systems. Among these attacks, a recent class of flatness-enhanced methods [3, 7, 8, 2] has emerged as a promising direction by linking the geometric flatness of the adversarial loss landscape to adversarial transferability. Flatness-enhanced methods craft adversarial examples with flatter neighborhood loss landscapes, enhancing cross-model robustness and achieving state-of-the-art transferability [8].
@@
Based on the empirical and theoretical analysis, we propose the Maximin Expected Flatness (MEF) attack to enhance adversarial transferability while reducing computational cost. Our key insight is to leverage the complementary strengths of average-case and worst-case flatness optimization, where average-case methods stabilize local exploitation while worst-case strategies preserve global exploration capability. MEF attack explicitly optimizes zeroth-order worst-neighborhood average flatness, a unified metric that internally computes neighborhood-level average-case flatness to stabilize local exploitation and externally identifies worst-case regions to guide global exploration. MEF implements this via gradient-guided neighborhood conditional sampling (NCS) for efficient worst-region identification and gradient balancing optimization (GBO) that reuses gradients across stages, reducing computation by 50% while enhancing transferability. By coupling these mechanisms, MEF achieves flattest loss landscape (as shown in Fig. 1) and enhanced adversarial transferability. Third-party evaluation on the widely-adopted TransferAttack benchmark shows MEF dominates 26 gradient-based methods as the new leader. Our main contributions are summarized as follows:
@@ II Related Work
This section systematically reviews adversarial machine learning literature through two lenses: (1) we survey transfer-based adversarial attacks, and (2) analyze various defense mechanisms, including both reactive and proactive defenses.
@@ II-A Adversarial Attacks
$\underset{\mathbf{x}^{adv}\in B_{\epsilon}(\mathbf{x})}{\max}J(\mathbf{x}^{adv},y;\mathcal{F}),$ (1)
@@
Data-driven Methods. Data-driven methods [14, 15, 17, 18] boost adversarial transferability by aggregating gradients over multiple augmented variants through data transformation strategies. The key distinction between these methods centers on the specific transformation operators. DI [14] optimizes over resized images, TI [15] applies spatial shifting, and SI [17] manipulates pixel scaling. Beyond these single-input spatial transformations, Admix [18] enhances diversity by blending cross-category images, while SSA [19] perturbs frequency domains to break spatial locality. To further amplify input variety, composite strategies integrate multiple transformations [20, 21]. However, such strategies incur significant computational overheads and induce semantic distortions, ultimately compromising their effectiveness.
@@
Model-driven Methods. Model-driven methods [22, 23, 24, 25] exploit geometric priors in surrogate models’ feature spaces to enhance adversarial transferability, primarily through two strategies: model ensemble aligns heterogeneous architectures via joint perturbation optimization to amplify spectral commonality with victims [24, 25], while surrogate refinement adversarially tunes surrogates or aligns gradients to reshape decision boundaries [22, 23]. However, model ensemble scales linearly with surrogates’ computational costs, and refinement struggles on large datasets due to complex regularization, limiting both to small-scale applications.
@@
Optimization-driven Methods. Optimization-driven methods [16, 7, 26, 27], enhance transferability by redesigning optimization objectives [26, 27] or algorithms [28, 16]. Advanced Objective strategies replace cross-entropy loss with feature discrepancy losses that align surrogate-victim model features [26, 27]. Generative Modeling strategies leverage generative networks to synthesize perturbations without iterative updates [28]. Gradient Stabilization strategies retain simple losses but refine gradient dynamics via momentum [16, 17, 4] or variance reduction [1]. Compared to the former two paradigms requiring complex losses or architectures, gradient stabilization strategies achieve transferability with minimal complexity, avoiding auxiliary data and models.
@@
Recent advances in gradient stabilization attacks have seen the emergence of flatness-enhanced methods, which propose that adversarial examples within flatter loss regions exhibit stronger transferability. These methods diverge into two strategies: those optimizing average-case flatness by aggregating gradients over local neighborhoods (e.g., FEM [3], APP [6]), and those pursuing worst-case flatness by minimizing gradient peaks (PGN [8]) or maximizing minimal loss (RAP [7]). However, existing works adopt fragmented flatness definitions while relying on empirically assumed correlations between flatness and transferability. Crucially, these methods typically employ flatness as a heuristic regularizer without a rigorous theoretical foundation linking specific geometric properties to the transferability gap. This theoretical void leads to suboptimal optimization directions. Our work bridges this gap by establishing unified flatness-transferability guarantees and adaptive optimization that harmonizes both flatness criteria.
@@ II-B Adversarial Defenses
Proactive Defenses. Proactive defenses aim to intrinsically harden model parameters against adversarial perturbations. The dominant approach, adversarial training (AT) [11, 29], reformulates robust model optimization as a min-max optimization problem, where the inner maximization generates perturbations to craft challenging adversarial examples, while the outer minimization enforces the model to correctly classify these perturbed samples. Variants like ensemble adversarial training (EAT) [30] augment robustness through multi-model perturbations during training, strengthening defense against cross-architecture transfer attacks. In contrast to AT’s empirical robustness, randomized smoothing [31] provides certifiable robustness by injecting Gaussian noise into inputs and certifying predictions over perturbed samples. While this method theoretically guarantees robustness within a certified radius, such radii are often smaller than adversarial perturbations defended by AT, rendering it ineffective against practical adversarial threats despite its theoretical appeal.
Reactive Defenses. Reactive defenses mitigate adversarial attacks without modifying model parameters. Adversarial purification [32, 33] projects inputs onto clean data manifolds using generative models [32, 33], while adversarial detection identifies adversarial examples via activation anomalies or confidence thresholds. These strategies offer plug-and-play deployment but suffer from two notable limitations: purification efficacy depends critically on generative model capacity [34, 35], and detectors can be bypassed by sophisticated adaptive attacks that mimic clean data statistics.
@@ III Limitations in Flatness Enhancement
The heuristic adoption of flatness concepts in flatness-enhanced methods lacks rigorous theoretical grounding while suffering from absence of systematic comparisons between flatness metrics, which constrains their effectiveness and efficiency. In this section, we first formalize flatness as a measurable geometric property, resolving definitional ambiguities across existing methods. Then we identify two inherent limitations in current flatness-enhanced methods. Finally, we distill the critical yet understudied challenges in balancing local exploitation and global exploration for real-world deployment.
@@ III-A K-th Order Flatness
    $\widehat{R}^{(n)}_{\xi}(\mathbf{x})\triangleq\max_{\mathbf{x}^{\prime}\in B_{\xi}(\mathbf{x})}\big\|\nabla^{n}J(\mathbf{x}^{\prime})-\nabla^{n}J(\mathbf{x})\big\|_{2},$ (2)
@@
    $\overline{R}^{(n)}_{\xi}(\mathbf{x})\triangleq\mathbb{E}_{\mathbf{x}^{\prime}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}))}\big[\big\|\nabla^{n}J(\mathbf{x}^{\prime})-\nabla^{n}J(\mathbf{x})\big\|_{2}\big],$ (3)
where $\mathbf{x}^{\prime}$ denotes a point in $\ell_{\infty}$-ball $B_{\xi}(\mathbf{x})=\big\{\mathbf{x}^{\prime}\in\mathbb{R}^{d}\,\big|\,\|\mathbf{x}^{\prime}-\mathbf{x}\|_{\infty}\leq\xi\big\}$ centered at $\mathbf{x}$ with radius $\xi$, $\nabla^{n}J(\mathbf{x})$ denotes the nn‑th order derivative of $J$ at $\mathbf{x}$ and $\mathrm{Unif}(B_{\xi}(\mathbf{x}))$ denotes the uniform distribution on the set $B_{\xi}(\mathbf{x})$.
[…]
@@ III-B Rethinking Flatness-Enhanced Methods
[…]
To validate the over-exploitation and over-exploration issues, we generate adversarial examples on Res-50 [9] using four flatness-enhanced methods (RAP [7], FEM [3], PGN [8], TPA [2]) for 100 iterations. We measure inter-update gradient cosine similarity during optimization and evaluate transfer attack success rates against Inc-v3 [37]. The gradient similarity serves as an indicator of the optimization dynamics: high similarity signifies consistent update directions for steady exploitation, whereas low values reflect frequent shifts in search directions for broad exploration.
The experimental results, illustrated in Fig. 2, reveal a clear empirical divergence. Worst-case flatness optimization exhibit significantly lower inter-iteration gradient similarities (mean similarity: RAP 0.10, PGN 0.25) compared to average-case counterparts (FEM 0.38, TPA 0.88) after convergence. The persistently low gradient similarity in worst-case optimization reveals an unstable oscillatory trajectory, an indicator of over-exploration driven by the hypersensitivity of the maximization objective to shifting local extrema in non-convex landscapes. In contrast, high similarity in average-case optimization suggests over-exploitation of local gradient consensus, leading to premature convergence on suboptimal plateaus as the expectation operator traps the search in dominant local directions. Ultimately, both pitfalls lead to premature stagnation of optimization, thereby strictly constraining the maximum achievable adversarial transferability.